the longest side of the triangular section is 8 feet shorter than twice the shorter side. the third side is 4 feet longer than the shortest side. the perimeter id 60 feet. how long is each side?
Answer by JulietG(1812)

A + B + C = 60
A = 2B - 8 (where A is the longest side, and B is the shortest side. We know C isn't the shortest side from the second sentence.)
C = B + 4
Substitute the values of A and C from the bottom equations into the first
A + B + C = 60
(2B-8) + B + (B+4) = 60
4B -4 = 60
Add 4 to each side
4B = 64
Divide each side by 4
B = 16
If the shortest side (B) is 16, then A is (2*16)-8, or 24; C is 16+4, or 20
A = 24
B = 16
C = 20